OpenStudy (anonymous):

What is the inverse of y = 3x + 9?

OpenStudy (anonymous):

-1/3x+9

OpenStudy (anonymous):

you guys mean y=9x+3 right?

OpenStudy (anonymous):

A. y= 1/3x-3 B. y=9x+3 C. y=1/3x+9 D. y=-1/3x-3

OpenStudy (anonymous):

still got y=3x+9 I think the actual answer is B

OpenStudy (anonymous):

cause were working with "inverse"

OpenStudy (anonymous):

Hi friend, you just replace x by y and y by x to get the inverse function. and then solve for y . to your problem, when replace you have x= 3y + 9 ---> 3y =x-9---> y = (x -9)/3 or y = 1/3x -3 . that is inverse function of your function hope this help

OpenStudy (anonymous):

so @Hoa our answer would be y=1/3x-3

OpenStudy (anonymous):

absolutely right. and that is the easiest way to get the inverse function

OpenStudy (anonymous):

thanks!